# Probabilistic Hash Embeddings for Online Learning of Categorical Features

### v1

@@ 1 Introduction
Categorical features occur in many high-value ML applications: finance (clements2020sequential), fraud detection (al2021financial), anomaly detection (han2022adbench), cybersecurity (sarker2020cybersecurity), medical diagnosis (shehab2022machine), recommendation systems (ko2022survey; lai2023adaembed) etc. A large vocabulary and embedding table are strong characteristics of these categorical feature-intensive applications. While assigning each item [^1] We use the phrase item to refer to a categorical value. its own row in a large embedding table typically improves accuracy, it comes at a cost. A larger embedding table requires more resources / memory to deploy and slows down execution.
@@
A common and well-established solution for learning categorical features at a large scale, without maintaining a long vocabulary and a large embedding table, is the use of hashing techniques (weinberger2009feature; tito2017hash; shi2020beyond; kang2021learning; lai2023adaembed; coleman2023unified). For any categorical value (typically in the form of strings), a fixed hash function maps it to a hashed value in a small, predetermined finite set. This mapped hashed value serves as the resulting feature value, indexing a row in a much smaller embedding table. These indexed embeddings, referred to as hash embeddings, are used in subsequent model training and inference.
@@
Hash collisions occur when different items share the same hashed value and embedding, potentially causing model performance drops. However, this issue can be mitigated in two ways: first by designing sophisticated operations of hashed values (weinberger2009feature); second by using multiple hash functions (tito2017hash; coleman2023unified). Large technology firms like Yahoo and Google have successfully incorporated this approach in their applications (weinberger2009feature; coleman2023unified).
@@
A major limitation of prior work is they focused on offline settings where data distributions are stationary and the entire test set is available in a batch fashion. In other words, the vocabulary of categorical items is fixed with no new or out-of-vocabulary items occurring during testing.
@@
Figure 1: On two tabular datasets, Mushroom and Adult, we split the data into groups based on a random partition of a categorical column’s vocabulary, such that each group has a disjoint vocabulary. We report the results before and after online learning on each group in the plots. The performance gaps motivate the need to learn representations of new items. Brackets are the columns used for splitting. Results are averaged on five runs. The partition detail is in Fig. 9.
@@
In many real applications, data arrives in a streaming fashion: (i) the vocabulary of categorical items can change, and/or (ii) the semantic meaning of an item can evolve. These characteristics present challenges to offline predictive models. Failure to adapt to the expanding vocabulary leads to a loss in predictive performance, as shown in Fig. 1 – in streaming binary classifications, new groups or sets of categorical items occur sequentially, and incrementally modeling these new items leads to significant accuracy improvement. This problem of expanding vocabulary is fairly common in practice: new products are added to a grocery store (cheng2023efficient), new usernames and application names appear in intrusion detection systems (siadati2017detecting; le2022rethinking), new patients arrive at a hospital, and so on.
@@
More severely, which is shown in our analysis and experiments below (Secs. 3.4 and 4), naively adapting to new items while using hashing techniques is subject to catastrophic forgetting (kirkpatrick2017overcoming). In hash embeddings, representations of two items may share parameters, updating one item’s embedding can adversely interfere with another, causing an effect like the model “forgets.” Furthermore, depending on the order of arrivals, the forgetting can be exacerbated. Consequently, hash embeddings are not yet fit for online learning in its vanilla form.
@@
In this paper, we employ Bayesian online learning to mitigate the forgetting issue. This approach is theoretically shown to be as effective as offline batch learning, regardless of the order in which items arrive (opper1999bayesian). To implement this, we model hash embeddings as stochastic and infer their posterior upon the arrival of new data. This approach mitigates the forgetting issue commonly encountered in the online updating of hash embeddings.
@@
Main Contributions: Our work proposes probabilistic hash embeddings (PHE) with Bayesian updates to handle dynamic vocabularies in an effective and efficient way. The intuition behind PHE stems from its benefits in (i) efficiency, as memory/number of model parameters is bounded and only a small number of parameters need to be updated online (ie., less forgetting, see experiments), and (ii) accuracy benefits since the Bayesian treatment provides an implicit regularization to trade-off forgetting and adaptation while maintaining invariance to item arrival order, without the need for specific dataset-dependent regularization design.
@@
We highlight PHE as a plug-in module, which can be applied to other probabilistic models such as Deep Kalman Filters (krishnan2015deep) and Neural Collaborative Filtering (he2017neural). The usage of PHE allows those models to handle unbounded items in their application areas in a principled way. We derive scalable variational inference algorithms to learn PHE and analyze why PHE is superior for online learning. We also derive results in explaining why PHE is superior for online learning. Empirically, our method outperforms baselines in three setups: sequential supervised learning with new items, conditional sequence modeling with increasing sequences and items, and a recommendation system with evolving user-item interactions.
@@
Organization: We survey related work in Sec. 2, present PHE, derive its inference algorithm in Sec. 3 and demonstrate PHE’s efficacy in Sec. 4 and conclude in Sec. 5. More details and limitation discussions are in supplementary materials.
@@ 2 Related Work
weinberger2009feature first proposed using hashing to handle unbounded number of categorical items. To improve on degradation due to hash collisions, serra2017getting used bloom filters. In recent times, tito2017hash; cheng2023efficient; coleman2023unified propose a shared embeddings across all categorical features for efficiency and using multiple hashing functions to reduce collisions. However, unlike our method, these are deterministic and are developed in offline learning settings. As shown in our experiments, deterministic hashing embeddings are vulnerable to evolving vocabularies.
Continual learning. Previous continual learning methods (kirkpatrick2017overcoming; nguyen2018variational; lopez2017gradient) were designed for continuous-valued features and do not address how to learn evolving categorical ones. This only existing approach, architecture-expanding methods (see expandabel embeddings EE in experiments) that dynamically expand the embedding tables, leads to unbounded memory usage (rusu2016progressive; yoon2017lifelong; jerfel2019reconciling). In contrast, PHE is the first to maintain constant memory while preserving accuracy and invariance to the order of category arrival.
Temporal and recommendation models. Temporal and recommendation models are important applications of categorical feature embeddings. One of our application models extends Deep Kalman Filters (DKF) (krishnan2015deep) to be applicable for evolving multi-task sequence modeling, while the original DKF assumes the vocabulary of categorical features is fixed. Similarly, previous recommendation methods (ko2022survey; shi2020beyond; kang2021learning; coleman2023unified) assume training data is given at once and the vocabulary of items is stationary.
Tabular data models. Categorical features exist in almost every tabular dataset. Handling tabular data requires handling categorical features as well. In online and continual learning settings, deep learning-based methods have been studied in recent years (huang2020tabtransformer; du2021tabularnet; liu2023incremental). However, all of these works assume that the vocabulary of categorical items are known and fixed up-front. Ours is the first online learning method, even for regular tabular data, that can handle increasing and unbounded items. While kimcarte use string embeddings from language models for open-vocabulary categorical/string-valued features in an offline setting, we focus on online settings. Moreover, ML applications cannot exploit representations from language models if strings are not semantically meaningful (like an anonymized ZIP code, IP address, etc.). We survey and discuss additional related work in Supp. D.
@@ 3 Methodology
Figure 2: Probabilistic hash embeddings for categorical feature $\mathbf{s}$}. For example, $\mathbf{s}$} can be usernames or anonymized strings. The whole module serves as $p(E_{{\mathbf{h}}_{\mathbf{s}}})$.
In this section, we first establish the necessary notations and introduce our proposed probabilistic hash embedding (PHE) module. Next, we derive an online inference algorithm for PHE. Finally, we analyze and explain why PHE is superior in online learning.
@@ 3.1 Notations and Problem Setup
We consider the problem of learning predictive ML models with categorical features in an online learning environment. We assume the model input and output dimensions are fixed but new categorical features may occur over time. [^2] We also assume categorical features are single-valued, that is, each feature is assigned with one category exactly. But our work is compatible with multi-valued features.
We denote categorical items by $s\in\mathcal{S}$. Typically, some features are of particular interest, and practitioners aim to predict these using ML models based on other features. We represent these features of interest, referred to as targets, by $y$. Depending on the task, $y$ can be either numeric or categorical. We use the subscript $i$ to index the $i$-th datum.
Hashing techniques use hash functions (e.g., MD5) to map categorical items to hashed values. These hashed values are then used to index rows of an embedding table. The range of hash values is usually much smaller than the vocabulary size of the original categorical items, thus achieving memory and execution efficiency. In this procedure, we let $h:\mathcal{S}\rightarrow\mathbb{N}_{<B}$ be a hash function that maps a string item to a hash value. The hash value has a pre-defined upper bound $B$, which is also known as the “bucket size”. For simplicity, we use $h_{s}$ to denote the hash value $h(s)$ of an item $s$. The hash value $h_{s}$ indexes one row in a hash embedding table $E\in\mathbb{R}^{B\times d}$, yielding the embedding representation $E_{h_{s}}$ of $s$. We use the same notation for both random variables and their sampled values where the meaning should be clear from the context.
@@ 3.2 Probabilistic Hash Embeddings (PHE)
To start off, we introduce our proposed encoding module for categorical items–probabilistic hash embeddings (PHE). A basic PHE has two components–a fixed hash function $h$ and a random hash embedding table $E$ with a prior distribution $p(E)$. In this work, we assume EE is Gaussian with independent entries. Given an item $s$, PHE looks up the $h_{s}$th row of $E$ as its embedding $E_{h_{s}}$, whose distribution is $p(E_{h_{s}})$.
A single hash function may result in two distinct inputs having the same hashing value, known as hash collisions, resulting in undistinguished hash embeddings. For size-$B$ buckets, the collision probability is proportional to $O(1/B)$. To further reduce the collision rate, we use universal hashing (carter1977universal). Namely, instead of utilizing one hash function, we use $K$ hash functions with different random seeds but the same range. Then the collision probability can be shown to reduce to $O(1/B^{K})$. Moreover, we keep the hash embedding table $E$ shared across $K$ hash functions, which keeps the model size bounded.
We now describe how PHE encodes an item $s$. This procedure is illustrated in Fig. 2. With $K$ hash functions, a categorical feature $s$ results in $K$ hash values ${\mathbf{h}}_{s}:=\{h^{(1)}_{s},\dots,h^{(K)}_{s}\}$, which in turn result in $K$ embeddings $\{E_{h^{(1)}_{s}},\dots,E_{h^{(K)}_{s}}\}$ where the $k$-th embedding $E_{h^{(k)}_{s}}$ is the looked-up embedding based on the $k$-th hash value $h^{(k)}_{s}$. The final representation $E_{{\mathbf{h}}_{s}}:=g(E_{h^{(1)}_{s}},\dots,E_{h^{(K)}_{s}})$ of $s$ is generated using an assemble function $g:{\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{d}$ to combine the $K$ embeddings. Typical choices of $g$ involve coordinate-wise summation, average, maximum, or minimum; other parametric choices of $g$ include weighted sums where the weights come from another parametric model. The output of PHE is a probabilistic embedding $E_{{\mathbf{h}}_{s}}$ with distribution $p(E_{{\mathbf{h}}_{s}})$. Thanks to the shared embedding table, the memory cost of PHE is $O(Bd)$, independent of the number of hash functions $K$.
One core ML task is to model correlations between two variables. A common query is to ask what the probability of observing a value of variable $y$ is given a categorical item $s$, namely $p(y|s)$. With PHE, we can compute it as
@@
where we follow the data generating process and treat ${\mathbf{h}}_{s}$ to be the same as $s$ (which holds in the absence of hash collisions). In computing Eq. 1, we can employ the Monte Carlo method to estimate the expectation. (Sec. G.1 discusses practical implementations.) In this way, one can answer probability queries conditioned on discrete features.
@@ 3.3 Online Learning of PHE
In this subsection, we derive scalable inference algorithms for PHE using a simple yet generic model. These algorithms can be adapted for other model variants. We first focus on the static setting before expanding to the online setting.
Given observations $\mathcal{D}:=\{(y_{i},{\mathbf{s}}_{i})\}_{i=1}^{N}$ where ${\mathbf{s}}_{i}$ represent a set of categorical features, we want to learn correlations between $y$ and $\mathbf{s}$} using PHE. We assume observations are conditionally independently and identically distributed (i.i.d.) (conditional on $E$) and have likelihood $p(y|E_{{\mathbf{h}}_{\mathbf{s}}})$. [^3] We use $E_{{\mathbf{h}}_{\mathbf{s}}}$ to denote the concatenation of hash embeddings for each item if $\mathbf{s}$} contains more than one categorical items. PHE places a prior $p(E)$ over the hash embedding table $E$ and infer the posterior given the observations $\mathcal{D}$. By Bayes rule, the posterior is $p(E|\mathcal{D})\propto p(E)p({\mathcal{D}}|E)=p(E)\prod_{i=1}^{N}p(y_{i}|E_{{\mathbf{h}}_{{\mathbf{s}}_{i}}})$, which is often intractable with complex likelihoods, except for a small subset of models.
Therefore, we turn to approximate inference and apply variational inference (blei2017variational; zhang2018advances) to learn an approximate posterior by minimizing the Kullback-Leibler (KL) divergence $D_{\mathrm{KL}}(q_{\lambda}(E)|p(E|\mathcal{D}))$ between a variational distribution and the true posterior. We assume the variational posterior factorizes as $q_{\lambda}(E)=\prod_{b=1}^{B}\prod_{j=1}^{d}q_{\lambda_{bj}}(E_{bj})$ and $q_{\lambda_{bj}}(E_{bj})$ takes a Gaussian form with parameters $\lambda_{bj}:=\{\mu_{bj}\in\mathbb{R},\sigma_{bj}\in\mathbb{R}\}$. Finding the optimal variational distribution that minimizes the KL divergence is equivalent to finding the corresponding optimal parameters $\lambda^{*}:=\{\lambda_{bj}^{*}\}$. In this paper, we also assume the prior has the same factorization as the variational posterior and each entry in the embedding table is independent and takes a standard Gaussian with zero mean and unit variance, ie., $p(E_{bj})=\mathcal{N}(0,1)$. Plugging $p(E)$ and $q_{\lambda}(E)$ into the KL divergence yields an objective function $\mathcal{L}(\lambda)$ (also known as the evidence lower bound, ELBO) to be maximized (see derivations in Supp. A):
@@
Suppose we observe a second dataset ${\mathcal{D}}_{1}:=\{(y_{i},{\mathbf{s}}_{i})\}_{i=1}^{N_{1}}$ to which we would like our model to adapt and still be effective to $\mathcal{D}$}. Bayes rule suggests a principal online learning iteration $p(E|{\mathcal{D}}_{1},{\mathcal{D}})\propto p(E|{\mathcal{D}})p({\mathcal{D}}_{1}|E)$ to accommodate both datasets without storing both. We iteratively replace the prior distribution with the previous posterior and repeat the inference procedure for the new posterior. This iteration assumes datasets $\mathcal{D}$ and $\mathcal{D}_{1}$ are conditionally i.i.d.
Variational inference produces an approximation of the true posterior. We thus use $q_{\lambda_{0}^{*}}(E)$ in place of $p(E|{\mathcal{D}})$ and infer $\tilde{p}(E|{\mathcal{D}}_{1},{\mathcal{D}})\propto q_{\lambda_{0}^{*}}(E)p({\mathcal{D}}_{1}|E)$. The inference may again face the same intractability issue as previous iterations, for which we resort to variational inference once more. We set a new variational distribution $q_{\lambda}(E)$ and a KL divergence $D_{\mathrm{KL}}(q_{\lambda}(E)|\tilde{p}(E|{\mathcal{D}}_{1},{\mathcal{D}}))$ to be minimized. Rewrite the KL divergence gives us a new ELBO (the objective function)
@@
Comparing to Eq. 2, the original prior $p(E)$ of EE is replaced with $q_{\lambda^{*}_{0}}(E)$. Upon optimization convergence, the new optimal variational distribution is an approximate posterior given both datasets ($\mathcal{D}$} and ${\mathcal{D}}_{1}$). When new datasets arrive in the future, we repeat the above online learning iteration to accommodate new datasets. Although this procedure bears resemblance to traditional continual learning (wang2023comprehensive; nguyen2018variational; li2021detecting), is different since we focus on changing discrete items. When an ML task is specified, we can readily plug in the specified likelihood model into Eqs. 2 and 3 and optimize.
@@
Most expressive models contain learnable parameters $\theta$ in their likelihood $p_{\theta}(y|E_{{\mathbf{h}}_{s}})$. One needs to learn $\theta$ in addition to inferring the posterior of $E$. Fortunately, a minor modification in our previous derived objective functions (Eqs. 2 and 3) can achieve this. We replace $p(y|E_{{\mathbf{h}}_{s}})$ with the parametric one $p_{\theta}(y|E_{{\mathbf{h}}_{s}})$ and maximize ${\mathcal{L}}(\lambda,\theta)$ with respect to $\{\lambda,\theta\}$. These new objective functions are viable and correspond to variational EM algorithms (rezende2014stochastic; kingma2013auto), for which we provide proofs in Sec. A.3. Let $\{\lambda^{*}_{0},\theta^{*}\}$ maximize the modified Eq. 2 ${\mathcal{L}}(\lambda,\theta)$. During online learning, we fix both $\theta^{*},\lambda_{0}^{*}$ in the modified objective function Eq. 3 ${\mathcal{L}}^{(1)}(\lambda;\lambda_{0}^{*},\theta^{*})$ for efficiently online updating the belief of hash embeddings (assuming changing categorical features).

### v2

@@ Abstract
Codehttps://github.com/aodongli/probabilistic-hash-embeddings
Extended versionhttps://arxiv.org/abs/2511.20893
@@ 1 Introduction
Categorical features occur in many high-value ML applications: finance (clements2020sequential), fraud detection (al2021financial), anomaly detection (han2022adbench), cybersecurity (sarker2020cybersecurity), medical diagnosis (shehab2022machine), recommendation systems (ko2022survey; lai2023adaembed) etc. Large vocabularies and embedding tables are strong characteristics of these categorical feature-intensive applications. While assigning each item [^1] We use the term item to refer to a categorical value. its own row in a large embedding table typically improves accuracy, this comes at a cost. Larger embedding tables require more resources and memory to deploy and slow down execution.
@@
A common and well-established solution for learning categorical features at scale, without maintaining a large vocabulary and embedding table, is the use of hashing techniques (weinberger2009feature; tito2017hash; shi2020beyond; kang2021learning; lai2023adaembed; coleman2023unified). For any categorical value (typically a string), a hash function maps it to an index in a small, fixed range. This index addresses a row in a much smaller embedding table. These rows, referred to as hash embeddings, are used in subsequent model training and inference.
@@
Hash collisions–when different items map to the same hashed value and thus share an embedding–can degrade model performance. This issue can be mitigated in two ways: first, by designing sophisticated operations on the hashed values (weinberger2009feature); second, by using multiple hash functions (tito2017hash; coleman2023unified). Large technology firms like Yahoo and Google have successfully incorporated this approach in their applications (weinberger2009feature; coleman2023unified).
@@
A major limitation of prior works is they focused on offline settings where data distributions are stationary and the entire test set is available in batch. In other words, the vocabulary of categorical items is fixed with no new or out-of-vocabulary items occurring during testing.
@@
Figure 1: On two tabular datasets, Mushroom and Adult, we split the data into groups based on random partitions of categorical column vocabularies, such that each group has disjoint vocabulary sets. We report the results before and after online learning on each group. The performance gaps motivate the need to learn representations of new items. Bracketed values indicate the splitting columns. Results are averaged over five runs. Partition details are in Fig. 9.
@@
In many real applications, data arrives in a streaming fashion: (i) the vocabulary of categorical items can change, and/or (ii) the meaning of an item can evolve. These characteristics present challenges to offline predictive models. Failure to adapt to the expanding vocabulary leads to a loss in predictive performance, as shown in Fig. 1–in streaming binary classifications, new groups or sets of categorical items occur sequentially, and modeling these new items incrementally leads to significant accuracy improvement. This problem of vocabulary expansion is fairly common in practice: new products are added to grocery stores (cheng2023efficient), new usernames and application names appear in intrusion detection systems (siadati2017detecting; le2022rethinking), new patients arrive at hospitals, and so on.
@@
More critically, as our analysis and experiments show (Secs. 3.4 and 4), naively adapting to new items while using hashing techniques suffers from catastrophic forgetting (kirkpatrick2017overcoming). In hash embeddings, different items may share embeddings, updating one can adversely affect another, causing the model to effectively “forget.” Furthermore, this forgetting can be exacerbated depending on the arrival order. Consequently, hash embeddings are not yet fit for online learning in their vanilla form.
@@
In this paper, we employ Bayesian online learning to mitigate the forgetting issue. This approach is theoretically shown to be as effective as offline batch learning, regardless of the order in which items arrive (opper1999bayesian). Specifically, we model hash embeddings as random variables and perform posterior inference upon observing new data. This probabilistic treatment enables continual adaptation to new items while preserving knowledge of previously seen ones.
@@
Main Contributions: Our work proposes probabilistic hash embeddings (PHE) with Bayesian updates to handle dynamic vocabularies effectively and efficiently. The intuition behind PHE stems from its benefits in (i) efficiency, as memory and the number of model parameters is bounded, with only a small number of parameters requiring online updates (ie., less forgetting), and (ii) accuracy, as the Bayesian treatment provides implicit regularization that balances forgetting and adaptation while maintaining invariance to item arrival order, without dataset-specific regularization design.
@@
We highlight PHE as a plug-in module, which can be applied to other probabilistic models such as Deep Kalman Filters (krishnan2015deep) and Neural Collaborative Filtering (he2017neural). The usage of PHE allows these models to handle unbounded items in their domains in a principled way. We derive scalable variational inference algorithms for learning PHE and provide theoretical analysis demonstrating PHE’s superiority for online learning. Empirically, our method outperforms baselines in three setups: online supervised learning with new items, conditional sequence modeling with growing vocabularies, and a recommendation system with evolving user-item interactions.
@@
Organization: We survey related work in Sec. 2, present PHE, derive its inference algorithms in Sec. 3, demonstrate PHE’s efficacy in Sec. 4, and conclude in Sec. 5. More details and limitation discussions are in supplementary materials.
@@ 2 Related Work
weinberger2009feature first proposed using hashing to handle an unbounded number of categorical items. To mitigate degradation due to hash collisions, serra2017getting used bloom filters. More recently, tito2017hash; cheng2023efficient; coleman2023unified proposes shared embeddings across all categorical features for efficiency and multiple hashing functions to reduce collisions. However, unlike our method, these approaches are deterministic and are developed in offline settings. As shown in our experiments, deterministic hashing embeddings are vulnerable to evolving vocabularies.
Continual learning.
Previous continual learning methods (kirkpatrick2017overcoming; nguyen2018variational; lopez2017gradient) were designed for continuous-valued features and do not address how to learn evolving categorical features. This only existing approach, architecture-expanding methods that dynamically expand the embedding tables (see EE in experiments), leads to unbounded memory usage (rusu2016progressive; yoon2017lifelong; jerfel2019reconciling). In contrast, PHE is the first to maintain constant memory while preserving accuracy and invariance to category arrival order.
Temporal and recommendation models.
Temporal and recommendation models are important applications for categorical feature embeddings. We extend Deep Kalman Filters (DKF) (krishnan2015deep) to handle evolving vocabularies, whereas the original DKF assumes a fixed vocabulary. Similarly, previous recommendation methods (ko2022survey; shi2020beyond; kang2021learning; coleman2023unified) assume training data is given in batch and the item vocabulary remains fixed.
Tabular data models.
Categorical features are ubiquitous in tabular datasets. In online and continual learning settings, deep learning-based methods have been studied in recent years (huang2020tabtransformer; du2021tabularnet; liu2023incremental). However, all of these works assume the vocabulary of categorical items is known and fixed in advance. Ours is the first online learning method for tabular data that can handle increasing and unbounded items. While kimcarte use string embeddings from language models for open-vocabulary categorical features in an offline setting, we focus on online settings. Moreover, language model representations cannot be exploited when strings lack semantic meaning (such as anonymized ZIP codes, IP addresses, etc.). We survey additional related work in Supp. D.
@@ 3 Methodology
Figure 2: PHE architecture. The input $\mathbf{s}$} can be usernames or anonymized strings. The entire module represents $p(E_{{\mathbf{h}}_{\mathbf{s}}})$.
We first establish the necessary notations and introduce our proposed probabilistic hash embedding (PHE) module. Next, we derive an online inference algorithm for PHE. Finally, we analyze and explain why PHE is superior in online learning.
@@ 3.1 Notations and Problem Setup
We consider the problem of learning predictive ML models with categorical features in an online learning environment. We assume the model input and output dimensions are fixed but new categorical features may occur over time. [^2] We also assume categorical features are single-valued, that is, each feature is assigned exactly one category. However, our work is compatible with multi-valued features.
We denote categorical items by $s\in\mathcal{S}$. Typically, practitioners aim to predict certain target features using ML models based on other input features. We denote these target features by $y$. Depending on the task, $y$ can be either numerical or categorical. We use subscript $i$ to denote the $i$-th data point.
Hashing techniques use hash functions (e.g., MD5) to map categorical items to hashed values. These hashed values are then used to index rows of an embedding table. The range of hash values is usually much smaller than the vocabulary size of the categorical items, thus achieving memory and computational efficiency. We define a hash function $h:\mathcal{S}\rightarrow\mathbb{N}_{<B}$ that maps an item to a hash value, where $B$ is a pre-defined upper bound known as the “bucket size.” For simplicity, we use $h_{s}$ to denote the hash value $h(s)$ of an item $s$. The hash value $h_{s}$ indexes a row in the hash embedding table $E\in\mathbb{R}^{B\times d}$, yielding the embedding representation $E_{h_{s}}$ of $s$. We use the same notation for both random variables and their realizations when the meaning is clear from context.
@@ 3.2 Probabilistic Hash Embeddings (PHE)
We first introduce our proposed encoding module for categorical items–probabilistic hash embeddings (PHE). A basic PHE has two components: a fixed hash function $h$ and a hash embedding table $E$ with a prior distribution $p(E)$. In this work, we assume each entry of EE is independently Gaussian-distributed. Given an item $s$, PHE looks up the $h_{s}$-th row of $E$ as its embedding $E_{h_{s}}$, whose distribution is $p(E_{h_{s}})$.
A single hash function may result in distinct inputs having the same hashing value, known as a hash collision, resulting in indistinguishable hash embeddings. With bucket size $B$, the collision probability is $O(1/B)$. To further reduce the collision rate, we employ universal hashing (carter1977universal). Namely, instead of using one hash function, we use $K$ hash functions with different random seeds but the same range. The collision probability then reduces to $O(1/B^{K})$. Moreover, we share the same hash embedding table $E$ across all $K$ hash functions, which keeps the model size bounded.
We now describe how PHE encodes an item $s$. This procedure is illustrated in Fig. 2. With $K$ hash functions, a categorical feature $s$ results in $K$ hash values ${\mathbf{h}}_{s}:=\{h^{(1)}_{s},\dots,h^{(K)}_{s}\}$, which index $K$ embeddings $\{E_{h^{(1)}_{s}},\dots,E_{h^{(K)}_{s}}\}$ where $E_{h^{(k)}_{s}}$ is the embedding retrieved using the $k$-th hash value $h^{(k)}_{s}$. The final representation $E_{{\mathbf{h}}_{s}}:=g(E_{h^{(1)}_{s}},\dots,E_{h^{(K)}_{s}})$ of $s$ is obtained using an assembly function $g:{\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{d}$ to combine the $K$ embeddings. Typical choices of $g$ involve coordinate-wise summation, average, maximum, or minimum; other parametric choices of $g$ include weighted sums where the weights come from another parametric model. The output of PHE is a probabilistic embedding $E_{{\mathbf{h}}_{s}}$ with distribution $p(E_{{\mathbf{h}}_{s}})$. Due to the shared embedding table, the memory cost of PHE is $O(Bd)$, independent of the number of hash functions $K$.
One core ML task is to model correlations between two variables. A common query is to compute the probability of observing a value of $y$ given a categorical item $s$, i.e., $p(y|s)$. With PHE, we can compute it as
@@
where we follow the data generating process and treat ${\mathbf{h}}_{s}$ to be the same as $s$ (which holds in the absence of hash collisions). To compute Eq. 1, we use Monte Carlo sampling to estimate the expectation. (See Sec. G.1 for practical implementation details.) This approach enables probabilistic inference conditioned on categorical features.
@@ 3.3 Online Learning of PHE
In this subsection, we derive scalable inference algorithms for PHE in a simple yet general model. These algorithms can be adapted to other model variants. We first focus on the static setting before expanding to the online setting.
Given observations $\mathcal{D}:=\{(y_{i},{\mathbf{s}}_{i})\}_{i=1}^{N}$ where ${\mathbf{s}}_{i}$ represents a set of categorical features, we want to learn correlations between $y$ and $\mathbf{s}$} using PHE. We assume observations are conditionally independently and identically distributed (i.i.d.) given $E$, with likelihood $p(y|E_{{\mathbf{h}}_{\mathbf{s}}})$. [^3] We use $E_{{\mathbf{h}}_{\mathbf{s}}}$ to denote the concatenation of hash embeddings for each item if $\mathbf{s}$} contains more than one categorical items. PHE places a prior $p(E)$ over the hash embedding table $E$ and infers the posterior given the observations $\mathcal{D}$. By Bayes rule, the posterior is $p(E|\mathcal{D})\propto p(E)p({\mathcal{D}}|E)=p(E)\prod_{i=1}^{N}p(y_{i}|E_{{\mathbf{h}}_{{\mathbf{s}}_{i}}})$, which is typically intractable with complex likelihoods.
Therefore, we turn to variational inference (blei2017variational; zhang2018advances) and learn an approximate posterior by minimizing the Kullback-Leibler (KL) divergence $D_{\mathrm{KL}}(q_{\lambda}(E)|p(E|\mathcal{D}))$ between a variational distribution and the true posterior. We assume the variational posterior factorizes as $q_{\lambda}(E)=\prod_{b=1}^{B}\prod_{j=1}^{d}q_{\lambda_{bj}}(E_{bj})$ where each factor $q_{\lambda_{bj}}(E_{bj})$ is Gaussian with parameters $\lambda_{bj}:=\{\mu_{bj}\in\mathbb{R},\sigma_{bj}\in\mathbb{R}\}$. Finding the optimal variational distribution that minimizes the KL divergence is equivalent to finding the optimal parameters $\lambda^{*}:=\{\lambda_{bj}^{*}\}$.
We assume the prior has the same factorization, with each entry independently following a standard Gaussian, i.e., $p(E_{bj})=\mathcal{N}(0,1)$. Substituting $p(E)$ and $q_{\lambda}(E)$ into the KL divergence yields an objective function $\mathcal{L}(\lambda)$ (also known as the evidence lower bound, ELBO) to be maximized (see derivations in Supp. A):
@@
Suppose we observe a second dataset ${\mathcal{D}}_{1}:=\{(y_{i},{\mathbf{s}}_{i})\}_{i=1}^{N_{1}}$ to which we would like our model to adapt while maintaining effectiveness on $\mathcal{D}$}. Bayes rule provides a principled online learning approach: $p(E|{\mathcal{D}}_{1},{\mathcal{D}})\propto p(E|{\mathcal{D}})p({\mathcal{D}}_{1}|E)$. This allows us to accommodate both datasets without storing them simultaneous. We iteratively replace the prior with the previous posterior and repeat the inference procedure. This iteration assumes $\mathcal{D}$ and $\mathcal{D}_{1}$ are conditionally i.i.d. given $E$.
Variational inference produces an approximation of the true posterior. We thus use $q_{\lambda_{0}^{*}}(E)$ in place of $p(E|{\mathcal{D}})$ and infer $\tilde{p}(E|{\mathcal{D}}_{1},{\mathcal{D}})\propto q_{\lambda_{0}^{*}}(E)p({\mathcal{D}}_{1}|E)$. This again faces intractability, so we apply variational inference with a new variational distribution $q_{\lambda}(E)$ to minimize $D_{\mathrm{KL}}(q_{\lambda}(E)|\tilde{p}(E|{\mathcal{D}}_{1},{\mathcal{D}}))$. Rewrite the KL divergence gives us a new ELBO (the objective function)
@@
Compared to Eq. 2, the original prior $p(E)$ of EE is replaced with the previous posterior approximation $q_{\lambda^{*}_{0}}(E)$. Upon optimization convergence, the new optimal variational distribution approximates the posterior given both datasets ($\mathcal{D}$} and ${\mathcal{D}}_{1}$). When new datasets arrive in the future, we repeat the above online learning iteration to accommodate new datasets. Although this procedure resembles traditional continual learning (wang2023comprehensive; nguyen2018variational; li2021detecting), it differs in focusing on evolving discrete vocabularies rather than continuous feature distributions. For any specified ML task, we can plug the corresponding likelihood model into Eqs. 2 and 3 and optimize.
@@
Most expressive models contain learnable parameters $\theta$ in their likelihood $p_{\theta}(y|E_{{\mathbf{h}}_{s}})$. One needs to learn $\theta$ in addition to inferring the posterior of $E$. Fortunately, a minor modification to our previously derived objective functions (Eqs. 2 and 3) can achieve this. We replace $p(y|E_{{\mathbf{h}}_{s}})$ with the parametric one $p_{\theta}(y|E_{{\mathbf{h}}_{s}})$ and maximize ${\mathcal{L}}(\lambda,\theta)$ with respect to $\{\lambda,\theta\}$ jointly. These new objective functions are viable and correspond to variational EM algorithms (rezende2014stochastic; kingma2013auto), for which we provide proofs in Sec. A.3. Let $\{\lambda^{*}_{0},\theta^{*}\}$ denote the optimizers of the modified Eq. 2, ${\mathcal{L}}(\lambda,\theta)$. During online learning, we fix both $\theta^{*}$ and $\lambda_{0}^{*}$ in the modified objective function Eq. 3, ${\mathcal{L}}^{(1)}(\lambda;\lambda_{0}^{*},\theta^{*})$, to efficiently update the posterior over hash embeddings as new categorical items arrive.